July 18 Day Shift Summary

Jason LaDue (jladue@cuhep.Colorado.EDU)
Sun, 18 Jul 1999 14:32:02 -0600 (MDT)

Shift Summary for owl shift, 07/18/99

On Shift: Jason LaDue, Maxim Bychov

Hours of beam delivered at all intensities: 6.5h
Hours of beam delivered at nominal intensity: 6.5h
Hours of E832 data (special runs not included): 5.75h
Hours of special runs (laser scans, muon runs...):0.0h

Accounting of lost time:

Changing tape and Stopping Runs.

========================================================================
Runs ended this shift:
Run 13941 18.1M BEAM_832
Run 13942 6.1M BEAM_832
Run 13943 6.0M BEAM_832

Tapes brought to FCC this shift:
QK*149-152, QKB052.

========================================================================
Accesses during shift:
None

========================================================================
Overview of shift:
Lose Beam for 2hours due to a Water Skid problem in the Linac.
CsI channel 1769 goes has a gain of ~15% for ~15 spills.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13941 367 07:24 Beam goes away, Ch13 says SY energy conservation.

07:37 Ch13 now says that Booster Water Skid problem.

07:51 Ch13 -> Linac Water Skid, no estimate

13941 08:15 Run stopped.

13942 08:28 Run started.

09:11 SDAQ button turns red. The scaler_check

09:22 Beam returns to Users

92 10:45 CsI channel 2856 has Lsr Stab RMS of 31.6% for 1
spill, then goes away.

107 10:51 KTEVANA errors, Unknown crate Stream 5. KFREND
fixes.

11:00 DAMP window is frozen, continue to run anyway.

13942 11:51 Run ends, the Tapes are full.

13943 12:17 Run begins.

14 12:33 CsI 1769 has a Gain Drift of -12.6%
Page Ed and Julie.

22 CsI 1769 Gain Drift now under the peak.
Add channel to the monitoring.

14:10 Blue Ice put around Bubbler.

90 14:29 CsI 1769 Gain Drift of -20.1% for one spill only.

106 CsI 1769 Gain Drift of -10%.
107 CsI 1769 Gain Drift of -20%, page Ed.

13943 14:45 End Run 13943 to allow for monitoring of 1769.

13944 14:56 Begin Run, first few spills show no problems
with CsI channel 1769

========================================================================
Follow up Notes on Previous Shift
Did not page Jason H. to replace AZLAT.

========================================================================
Items to Follow up on Next Shift

Page Jason H. if there is a significant amount of down time.
CsI channel 1769 had a bad gain for a few spills.